### v1

# Divergent Ensemble Networks: Enhancing Uncertainty Estimation with Shared Representations and Independent Branching

Abstract
Ensemble learning has proven effective in improving predictive performance and estimating uncertainty in neural networks. However, conventional ensemble methods often suffer from redundant parameter usage and computational inefficiencies due to entirely independent network training. To address these challenges, we propose the Divergent Ensemble Network (DEN), a novel architecture that combines shared representation learning with independent branching. DEN employs a shared input layer to capture common features across all branches, followed by divergent, independently trainable layers that form an ensemble. This shared-to-branching structure reduces parameter redundancy while maintaining ensemble diversity, enabling efficient and scalable learning.
Keywords:
uncertainty estimation, deep learning, artificial intelligence
1 Introduction
In controlled conditions and simulations, contemporary deep-learning networks have shown remarkable performance. However, when confronted with unexplored data, unforeseen fluctuations, or noisy inputs, its dependability frequently declines in real-world circumstances, potentially resulting in predictions that are not accurate. This raises concerns about the dependability of these networks in practical applications. [6]. To address this challenge, leveraging uncertainty estimation was proposed as a solution to enhance their robustness and ensure real-world applicability. Neural networks struggle with quantifying predictive uncertainty and often generate overly confident predictions. Such overconfident incorrect predictions can lead to detrimental results[8].
In deep learning, there are two kinds of uncertainties: aleatoric uncertainty, which originates from the inherent noise or randomness in the data and cannot be diminished by altering the model, while epistemic uncertainty comes from a lack of knowledge or information regarding the model, and can be decreased by enhancing the model design or gathering additional data [5]. For aleatoric uncertainty, the entropy of the predictive distribution is used to quantify the uncertainty [13]. Epistemic uncertainty is quantified through model variability. Several models have been proposed to address these types of uncertainty. Bayesian Neural Networks (BNNs) model uncertainty by treating network weights as trainable parameters [10]. Monte Carlo Dropout approximates the posterior distribution using dropout during inference [3]. Ensemble methods, such as Bagging, reduce uncertainty by averaging predictions from multiple models [1]. Bootstrapping is also used for uncertainty estimation which involves training multiple models on resampled subsets of the data [7]. Deep Ensembles combine multiple models to estimate uncertainty by evaluating their predictions’ variance [4].
The biggest challenge in the ensemble is to reduce redundancy while preserving its diversity. While techniques like Monte Carlo (MC) Dropout approximate ensembles by stochastic regularization, but they may compromise the independence of predictions [14]. Similarly, approaches like Batch Ensemble aim to reduce computation but struggle with flexibility in capturing model variance. This trade-off between computational efficiency and one of the primary obstacles to ensemble learning is still ensemble diversity. To address these challenges, we propose the Divergent Ensemble Network (DEN), a novel architecture that combines the benefits of shared representation learning and ensemble diversity. DEN features a shared input layer that captures a common representation across all ensemble members, followed by independent branching layers that train separately to maintain prediction diversity. This architecture balances computational efficiency with predictive robustness, making it suitable for scenarios where both accuracy and uncertainty estimation are critical.
In this study, we assess DEN across various tasks, emphasizing its ability to estimate uncertainty while preserving predictive accuracy. Our findings indicate that DEN surpasses conventional ensembles in computational efficiency and achieves similar or superior uncertainty estimation. Additionally, its scalable nature allows for adaptability in a wide range of applications, including those that require real-time predictions. The remainder of this paper is structured as follows: Section 2 provides an overview of the DEN architecture and its training approach. Section 3 showcases the findings, while Section 4 explores the applications and constraints, and Section 5 offers concluding remarks. The code is available on GitHub https://github.com/Arker123/Divergent-Ensemble-Networks.
2 Architecture
2.1 Problem setup and High-level summary
We assume that the training dataset $\mathcal{D}$ consists of $N$ Independent and Identically Distributed (i.i.d) data points. For classification problems, the label is assumed to be one of $K$ classes, that is $y\in\{1,\dots,K\}$. For regression problems, the label is assumed to be real-valued, that is $y\in\mathbb{R}$.
Given the input features $x$, we use a neural network to model the probabilistic predictive distribution $p_{\theta}(y|x)$ over the labels, where $\theta$ are the parameters of the neural network.
The Divergent Ensemble Network (DEN) uses a shared-to-branching architecture for ensemble learning. A common input layer processes shared features, followed by a single shared representation layer that extracts fundamental features. After this, the network diverges into multiple independent branches, each trained with distinct parameters to have independent predictions. These branches operate on the shared features and produce independent outputs. This model thus provides independent predictions while also giving faster predictions.
Figure 1: Visual representation of the proposed neural network architecture. The number of branches and the nodes in each branch are hyper-parameters and varies from problem to problem.
The shared layer in DEN ensures efficient parameter usage by processing input features once, reducing redundancy. Meanwhile, the divergent branches maintain the independence necessary for effective ensemble learning. By decoupling shared representation from individual predictions, DEN achieves a unique balance between computational efficiency and predictive variance.
2.2 Shared Input and Common Representation Layer
A shared input layer processes the raw input features at the beginning of the DEN architecture. By capturing shared representations among all ensemble members, this layer significantly reduces computational redundancy. By utilizing a shared feature extractor, DEN allows the best possible use of parameters while retaining high-quality input representations for all downstream branches. A fully connected layer with neurons makes up the shared input layer; the number of neurons is a hyperparameter which is application dependent.
2.3 Divergent Branches for Independent Predictions
The architecture divides into several independent networks after the shared layers. Each branch processes the shared representation further, using distinct weights and biases to enable independent predictions. These branches act as independent ensemble members while benefiting from the common input processing. Each branch consists of multiple fully connected layers with the activation functions also among the hyperparameter, and dropout layers in case improves the accuracy.
2.4 Loss Function and Optimization
Each branch is trained independently using a loss function specific to the task and each branch gets to see the data in a loop so that the common layers are trained. For classification tasks, the softmax cross-entropy loss is typically used, and mean squared error for the regression tasks. Additionally, ensemble diversity is encouraged by training branches independently, ensuring variance in predictions and robust uncertainty estimation.
3 Observations and Results
Firstly, we used the MNIST dataset to train our model. The MNIST dataset is a large, well-known database of handwritten digits comprising 70,000 images, of which 60,000 are used for training and 10,000 for testing. Each image is a grayscale representation of a digit (0–9) with a resolution of 28×28 pixels. As one of the benchmark datasets in machine learning and computer vision, MNIST provides a simple yet effective framework for evaluating classification algorithms.
To further test our model’s robustness, we utilized the NotMNIST [^1] Available at http://yaroslavvb.blogspot.co.uk/2011/09/notmnist-dataset.html dataset, which offers a similar structure but focuses on the classification of letters A through J in a variety of fonts. The dataset comprises 18,724 training images and 1,872 testing images, all grayscale and sized at 28×28 pixels.
Figure 2: Comparison of the MNIST and NotMNIST Datasets.
After comparing the results with the MNIST dataset, we extended our evaluation to test the model’s ability to quantify uncertainty by using it on the NotMNIST datasets. Then the NotMNIST dataset contains not only numbers but also letters, which provides a way to test how it performs on unseen data. By analyzing the model’s performance on these NotMNIST datasets, we aimed to evaluate how effectively it distinguishes between familiar (in-distribution) and unfamiliar (out-of-distribution) data. This approach has been widely practiced [8] involving unpredictable or previously unseen data.
We ran the model on an Intel i5 12th gen processor, utilizing only a single core for performance comparison purposes. We found an  6x improvement in execution time compared to other models Table 1 , making it suitable for real-time predictions. We tested the model on 9000 examples from each class, and evaluated the uncertainty on out-of-distribution examples from unseen classes. Moreover, overconfident predictions on unseen classes pose a challenge for reliable deployment of deep learning models in real world applications, hence, the predictions to exhibit higher uncertainty for the test data being much different from training data is expected. To test if the proposed method possesses this desirable property, we train a Multilayer perceptron (MLP) on the standard MNIST train/test split using the proposed neural architecture. However, in addition to the regular test set with known classes, we also evaluated it on a test set containing unknown classes. For this we used the test split of the NotMNIST dataset where the images have the same size as in the MNIST dataset, however, the labels are alphabets instead of digits. Though, we do not have access to the true conditional probabilities, we expect the predictions to be closer to a uniform value on unseen classes compared to the known classes where the predictive probabilities should concentrate on the true targets. We evaluated the entropy of the predictive distribution and used this to evaluate the quality of the uncertainty estimates. The results are shown in Table 1 and Table 2
Table 1: Evaluation Results (MNIST)
Model Ensemble Accuracy (%) Single Model Accuracy (%) Average Inference Time (s)
Ensemble 98.56 97.78 0.263453
MC Dropout 97.33 97.67 0.158070
Bootstrap 97.67 96.67 0.277438
Model DEN 98.78 98.44 0.056009
Table 2: Evaluation Results (NotMNIST)
Run Accuracy Average Confidence Average Variance Average Entropy Inference Time (s)
Ensemble 10.22 0.6754 0.0406 0.6707 0.228371
MC Dropout 9.78 0.7545 0.0053 0.5987 0.164834
Bootstrap 10.78 0.6691 0.0416 0.6853 0.229866
Model DEN 10.78 0.7019 0.0375 0.6125 0.045510
For the MNIST dataset, our model, outperforms the existing Ensemble, MC Dropout and Bootstrap methods by about 6x, and about 4x in case of NotMNIST dataset. All of the methods have low entropy as expected. MC-dropout seems to give high confidence predictions for some of the test examples, even for unseen classes. Such overconfident wrong predictions can be problematic in practice when tested on a mixture of known and unknown classes. DEN, Bootstrap and Ensemble produces higher uncertainty on unseen classes.
4 Applications
Uncertainty estimation methods have demonstrated significant potential in various domains of computer vision and medical image analysis [11]. In the field of Computer Vision and Image Processing, deep ensembles are particularly effective for detecting epistemic uncertainties, such as out-of-distribution (OoD) issues. They are also crucial in adversarial sample detection, where they help identify and mitigate vulnerabilities in neural networks [11]. Furthermore, the integration of ensemble methods with probabilistic embeddings has shown promise in tasks that require quantified uncertainties [12]. In the domain of continual learning (CL), deep ensembles have been utilized to address uncertainty, especially in variational continual learning methods, thus enabling robust learning in dynamic environments[15].
In Medical Image Analysis, ensembles find applications in confidence calibration and uncertainty estimation[9]. It can also be used for the estimation of uncertainties in robots [2]
5 Conclusion
This model achieved significant results on standard datasets for uncertainty estimation. The proposed architecture requires less space and computation time compared to traditional ensemble architectures. While Monte Carlo methods are more space-efficient than the proposed model, Gaussian Neural Networks, ensembles, and Monte Carlo approaches require multiple inferences, increasing the time needed to generate results. In contrast, the proposed model is well-suited for time-sensitive and real-time applications.